Compute sales for target net income.

(LO 5), AP For Flynn Company, variable costs are 70% of sales, and fixed costs are $195,000. Management's net income goal is $75,000. Compute the required sales in dollars needed to achieve management's target net income of $75,000. (Use the contribution margin approach.)
Answer #1

Contribution margin =Sales-Variable costs

=(100-70)=30%

Target Contribution margin=Fixed cost+Target net income

=(195,000+75,000)=$270,000

Hence required sales=Target Contribution margin/Contribution margin ratio

=$270,000/0.3

=$900,000.